a man can row downstream 32km in 6 hours.he can row 14km in same time if it is upstream,then the velocity of the stream is?
1
Expert's answer
2012-12-11T07:27:10-0500
The average velocity of man rowing downstream is& 32/6 km/h
The average velocity of man rowing upstream is& 14/6 km/h
The difference between average velocities is caused by the stream velocity. So the stream velocity can be found in such way:
Let the stream velocity& (Vs)=x, then
when man rows downstream he has his own velocity (Vm) and the stream velocity
Vmd=Vm+Vs
When man rows upstream he has his own velocity (Vm) and the stream velocity, which creates an obstacle for moving.
Vmu=Vm-Vs.
So, Vmd=Vm+Vs &
32/6=Vm+x => Vm=32/6-x
& Vmu=Vm-Vs
14/6=Vm-x => Vm=14/6+6

Hence 32/6-x=14/6+x => x=3/2 km/h
the velocity of the stream is 3/2 km/h